Show that if π‘Žπ‘‘βˆ’π‘π‘ β‰  0, then the function 𝑓(π‘₯)=π‘Žπ‘₯+𝑏𝑐π‘₯+𝑑 is one-to-one and find its inverse.


1
Expert's answer
2021-10-18T13:19:34-0400

If ad-bc=0 then:

Β a/c = b/d

y = a/c = b/d

So, y is a constant and is not one-to-one.


Finding the inverse function:

Let y = f(x) = (ax+b) / (cx+d)

So y (cx+d) = ax+b

So cxy + dy = ax+b

So cxy - ax = b - dy

So x (cy - a) = b - dy

So x = (b - dy) / (cy - a)

But x = f-1(y)

So f-1(y) = (b - dy) / (cy - a)

So f-1(x) = (b - dx) / (cx - a)


The fact that we have managed to find the inverse function means that f is one-one.